A 96% confidence interval for GPA values is (2.67, 3.15). What is the estimated mean GPA of the population, and what is the margin of error?
Answer #1

Solution :

Confidence Interval is Mean \pm ME

For 96% confidence Interval, the value of Z is 2.054

Upper Limit is 3.15

Lower Limit is 2.67

UL = 1 + 2x = 3.15 ........................(1)

LL = - 2x 2.67......................(2)

Adding (1) and (2), we get

2x = 2.6743.15 = 5.82

167 = zgs=1t:

b) Margin of error is

M+2 X = 3.15

2.91 + 2X == 3.15 Vn

= 3.15 – 2.91 = 0.24

ME = 0.24